OpenStudy (anonymous):

Need answers to this question so that i can compare your solutions to mine. This is a question on Satellite orbit A satellite with m = 450 kg should be placed in a circular orbit 430 km above earth’s surface. a) What speed and period must the satellite have? b) How much work has to be done to place the satellite in orbit? (Neglect air resistance and earth rotation) Please use: mE = 5.98 x10^24 kg, rE = 6.38 x10^6 m @Physics

OpenStudy (anonymous):

the first part of the question GmM/r^2=F=ma since its a circular orbit F will be mv^2/r so (mv^2)/r=GmM/r^2 for period omega w = 2pi/T T is the period and w is also equal to v/r so T=2pi/w henceforth T=2pi/(v/r)=2pi.r/v

OpenStudy (anonymous):

2nd part is tricky for circulat motion the work should be zero the motion is periodic and the total displacement is zero so the work is zero

OpenStudy (jamesj):

The second part is asking what is the work to GET the satellite in orbit; not what the work is when it is in orbit.

OpenStudy (anonymous):

my bad you re correct james